find the value of one edge if the volume of a cube is 6 times the area of one side
Answer by lwsshak3(11628)

find the value of one edge if the volume of a cube is 6 times the area of one side
**
let x=one edge of a cube
area of one side=x^2
volume of cube=x^3
x^3=6x^2
x^3/x^2=6
x=6
edge of cube=6